# TACLR: A Scalable and Efficient Retrieval-based Method for Industrial Product Attribute Value Identification

### v2

@@ 4.2 Metrics
Since most attributes in the taxonomy are exclusive (i.e., each product can have at most one value per attribute), we evaluate PAVI methods using micro-averaged Precision@1, Recall@1, and F1@1 scores.
For each attribute, the ground truth is a set of values $\mathcal{V}$ from the taxonomy. If the ground truth is empty ($\varnothing$), a correct prediction (True Negative, TN) occurs when the model also predicts an empty set; otherwise, it is a False Positive (FP). When the ground truth is not empty, the model’s top-1 output is a True Positive (TP) if it matches any ground truth value. Predicting an empty set in this case results in a False Negative (FN), while mismatched predictions are both False Positives (FP) and False Negatives (FN), as it simultaneously introduces an error and misses the correct value. [^3] In prior work Shinzato et al. (2023), metrics did not account for the FP case, and FP & FN cases were counted as FP only. We adopt more stringent metrics. Table 3 summarizes these outcomes. Final precision, recall, and F1 scores are computed by aggregating TP, FP, and FN counts across the dataset for a comprehensive performance evaluation.
Table 3: Confusion matrix comparing labeled value set with predicted value and their corresponding outcomes.
@@
Table 4: Performance comparison of different methods on Ecom-PAVI and WDC-PAVE. "F1 Excl." refers to the F1 score calculated without measurement attributes (e.g., width and height), which require unit conversion reasoning.
Paradigm Method Ecom-PAVI WDC-PAVE
Precision Recall F1 Score Precision Recall F1 Score F1 Excl.
@@ 4.3 Baselines
We evaluate our retrieval-based method TACLR against classification and generation baselines. [^4] Extraction-based baselines, such as NER or QA models, are widely used for PAVE, but they are excluded from our comparison due to the lack of a standard normalization method, which makes fair evaluation challenging. For implementation details, refer to Appendix A.
BERT-CLS. This baseline frames PAVI as a multi-label classification task, treating each category-attribute-value tuple as an independent label. The model is fine-tuned to predict matches, with label masking applied to exclude irrelevant labels for each category, following Chen et al. (2022). The model outputs a probability distribution over values and selects the highest probability value for each attribute. If no probability exceeds a specified threshold, the prediction is set to be empty.
LLMs. For generation-based baselines, we utilize state-of-the-art open-source LLMs, including Llama3.1-7B Llama Team (2024) and Qwen2.5-7B Qwen Team (2024). These models are initially evaluated in zero-shot and few-shot settings using a template adapted from Brinkmann et al. (2024), which incorporates the category, attribute, and product profile along with detailed value normalization guidelines. We also fine-tune the LLMs on task-specific data to predict attribute values in JSON format. A greedy decoding strategy is applied to ensure reproducibility.
@@ 5 Results
Table 5: Inference efficiency comparison on Ecom-PAVI (Throughput in samples/second).
@@
5.1 Main Results
Table 4 presents the performance comparison between our retrieval-based method TACLR and classification- and generation-based baselines on Ecom-PAVI and WDC-PAVE. On Ecom-PAVI, TACLR achieves the highest F1 score of 86.2%, surpassing the fine-tuned Llama3.1, which obtains an F1 score of 84.7%. Notably, TACLR excels in recall, achieving 87.1% compared to Llama3.1’s 82.7%. On WDC-PAVE, TACLR achieves the highest F1 Excl. score of 80.3%, which excludes measurement attributes requiring reasoning ability for unit normalization. This result highlights TACLR’s effectiveness and robustness in addressing general PAVI across diverse datasets.
The classification-based method, BERT-CLS, shows the weakest performance on both datasets. It achieves an F1 score of 50.5% on Ecom-PAVI, but its performance drops drastically on WDC-PAVE, where it only attains an F1 score of 20.5%. This underscores the limitations of classification-based approaches in generalization, including their inability to adapt to OOD values.
Among the generation-based methods, few-shot and fine-tuning consistently improve performance over zero-shot settings. For example, on Ecom-PAVI, Llama3.1 achieves F1 scores of 35.7%, 38.6%, and 84.7% in zero-shot, few-shot, and fine-tuned settings, respectively. Similarly, Qwen2.5 achieves F1 scores of 48.4%, 51.4%, and 81.7% in the corresponding settings. On WDC-PAVE, however, fine-tuned LLMs exhibit weaker generalization due to the scarcity of training data. Few-shot learning proves more robust in this dataset, achieving an F1 score of 75.0% for Llama3.1 and 72.2% for Qwen2.5.
@@ 5.1 Main Results
(c) Comparison on data domains.
@@
5.2 Inference Efficiency
Table 5 compares the inference efficiency of PAVI methods under identical evaluation conditions. We employed a naive PyTorch implementation without speed optimizations and used the largest batch size that avoids out-of-memory errors. All experiments were conducted on a machine equipped with one NVIDIA V100 GPU.
TACLR achieves a strong balance between performance and efficiency, with an inference time of 12.7 ms and a throughput of 630 samples per second. In contrast, generation-based methods, such as Llama3.1 (few-shot) and Qwen2.5 (few-shot), exhibit significantly longer inference times (124.8 ms and 98.4 ms, respectively) and lower throughputs (64 and 81 samples per second), highlighting the computational overhead of LLMs in high-load scenarios. While BERT-CLS delivers the fastest inference time and highest throughput, its inability to handle OOD values and limited capacity restrict its effectiveness in practical applications.
5.3 Analysis
@@
Impact of Taxonomy-Aware Negative Sampling. Figure 4 compares the proposed taxonomy-aware sampling (§3.3) with in-batch sampling across different sample sizes. As the number of sampled values increases, F1 score consistently improve, corroborating findings from Chen et al. (2020). With in-batch sampling as the baseline, the model achieves an F1 score of 53.3% with a sample size of 128. In contrast, taxonomy-aware sampling significantly outperforms this baseline, with an F1 score improving from 84.0% to 86.2% as the sample size increases from 16 to 128. These results highlight the superiority of taxonomy-aware sampling, which leverages the structure of attribute taxonomy to generate more challenging negative examples, enhancing the model’s recognition capabilities.
Comparison of Dynamic and Static Thresholds. Figure 5(a) evaluates the dynamic, learnable thresholds (§3.4) against the best static thresholds of 0.6, 0.65, and 0.7, which were chosen based on their performance on the validation set. The dynamic threshold achieves the highest F1 score of 86.2%, surpassing the static thresholds, which yield F1 scores of 75.5%, 80.2%, and 78.2%, respectively. Static thresholds exhibit a clear trend: as the threshold increases, precision rises (from 65.1% to 84.5%) while recall diminishes (from 90.0% to 72.8%). In contrast, the dynamic threshold balances precision (85.4%) and recall (87.1%) effectively, eliminating the need for extensive hyperparameter tuning across category-attribute pairs. This adaptability makes dynamic thresholds a practical choice.
@@
Performance Gains from Context-Rich Prompts. The influence of varying value prompt templates on the PAVI task is shown in Figure 5(b). Using only the value as a prompt achieves an F1 score of 83.2%. Adding category information raises the F1 score to 83.9%, while incorporating attribute information further improves it to 85.4%. The most comprehensive template, combining category, attribute, and value information (e.g., "A {category} with {attribute} being {value}"), achieves the highest F1 score of 86.2%. These results are consistent with prior work Radford et al. (2021), highlighting that context-rich prompts enhance the model’s discriminative performance.
Zero-Shot Generalization Across Data Domains. Figure 5(c) presents the results of zero-shot transfer experiments, evaluating the model’s generalization across unseen categories and values. The in-domain split achieves an F1 score of 88.7%, while performance decreases for cross-category and cross-value splits, which attain F1 scores of 80.2% and 78.2%, respectively. These declines reflect the inherent challenges of adapting to dynamic attribute taxonomies in OOD domains. Nevertheless, the overall F1 score of 86.2% demonstrates the robust generalization capabilities of TACLR.
@@ 6 Conclusion
In this work, we present TACLR, a novel approach for retrieval-based PAVI. By formulating PAVI as an information retrieval problem, TACLR enables the inference of implicit values, generalization to OOD values, and the production of normalized outputs. Building on this framework, TACLR employs contrastive training with taxonomy-aware sampling and adaptive inference with dynamic thresholds to enhance retrieval performance and scalability.
Comprehensive experiments on proprietary and public datasets demonstrated TACLR’s superiority over classification- and generation-based baselines. Notably, TACLR achieved an F1 score of 86.2% on the large-scale Ecom-PAVI dataset. Efficiency analysis further highlighted its advantage, achieving significantly faster inference speeds than generation-based methods. Beyond these experimental results, TACLR has been successfully deployed on a real-world e-commerce platform, processing millions of product listings daily and seamlessly adapting to dynamic attribute taxonomies, making it a practical solution for large-scale industrial applications.
@@ 7 Limitations
While TACLR demonstrates effectiveness in processing textual product profiles, it does not currently leverage multimodal information, such as images or videos. Multimodal data could provide valuable complementary context for attributes that are challenging to infer from text alone (e.g., visual attributes like color or texture). Incorporating multimodal capabilities may further enhance the model’s ability to identify attribute values with greater accuracy and comprehensiveness.
@@ Appendix A Implementation Details
We utilize pre-trained RoBERTa-base models Liu et al. (2019); Cui et al. (2020), augmented with a linear projection layer, to encode both the item profile and the value prompt. The embedding dimension is set to 256. For each product-attribute pair, we sample up to 128 values, which include a null value, an optional positive value (no positive value is selected when absent for the product-attribute), and negative values sampled from the same category-attribute pair. The temperature parameter for contrastive learning is set to 0.05. The models is fine-tuned using the AdamW optimizer with a batch size of 32 and a learning rate of 2e-5, over a maximum of 5 epochs. Hyperparameters and the best model checkpoints are selected based on the F1 score on the validation set.
@@ Appendix B Deployment
The proposed TACLR has been successfully integrated into key functionalities of an e-commerce platform, including product listing, search, recommendation, and price estimation. The system is designed to be highly scalable and efficiently process millions of products daily.
In the product listing process, TACLR automatically identifies attribute-value pairs from user-provided titles and descriptions, significantly reducing manual effort and errors while improving the quality of structured information.
For product search, improved structured information directly enhances lexical retrieval, leading to more accurate matching with user queries. Meanwhile, it enriches product features, consequently enhancing personalized recommendations.
In the area of price estimation, TACLR identifies key attributes that influence pricing, leading to more accurate price predictions. This provides both sellers and buyers with reliable, market-aligned information.

### v4

@@ 4.2 Metrics
Since most attributes in the taxonomy are exclusive, i.e., each product can have at most one value per attribute, we evaluate PAVI methods using micro-averaged precision@1, recall@1, and F1 score@1.
For each attribute, the ground truth is a set of values $\mathcal{V}$ from the taxonomy. If the ground truth set is empty ($\varnothing$), a correct prediction (True Negative, TN) occurs when the model also predicts an empty set; otherwise, it is a False Positive (FP). When the ground truth set is not empty, the model’s top-1 output is a True Positive (TP) if it matches any ground truth value. Predicting an empty set in this case results in a False Negative (FN), while mismatched predictions are both False Positives (FP) and False Negatives (FN), as it simultaneously introduces an error and misses the correct value. Table 3 summarizes these outcomes [^6] In prior work Shinzato et al. (2023), metrics did not account for the FP case, and FP & FN cases were counted as FP only. We adopt more stringent metrics in this paper.. Final precision, recall, and F1 scores are computed by aggregating TP, FP, and FN counts across the dataset, providing a comprehensive performance evaluation.
Table 3: Confusion matrix comparing ground truth value set with predicted top-1 value.
@@
Table 4: Performance comparison of classification-, generation-, and retrieval-based methods on Xianyu-PAVI and WDC-PAVE. "F1 Excl." denotes the F1 score computed excluding measurement attributes (e.g., width and height), which require complex unit normalization reasoning.
Paradigm Method Xianyu-PAVI WDC-PAVE
Precision Recall F1 Precision Recall F1 F1 Excl.
@@ 4.3 Baselines
We evaluate our retrieval-based method, TACLR, against classification and generation baselines [^7] Extraction baselines, such as NER and QA models, are not included because (1) they produce unnormalized outputs, requiring an established normalization strategy for fair comparison, and (2) the large-scale Xianyu-PAVI lacks token-level annotations (e.g., BIO tags) necessary for these methods.. For implementation details, refer to Appendix A.
BERT-CLS. This baseline frames PAVI as a multi-label classification task, treating each category-attribute-value tuple as an independent label. The model is fine-tuned to predict matches, with label masking applied to exclude irrelevant labels for each category, following Chen et al. (2022). The model outputs a probability distribution over values and selects the highest-probability value for each attribute. If no probability exceeds a specified threshold, the prediction is set to empty.
LLM. For generation-based baselines, we utilize state-of-the-art open-source LLMs, including Llama3.1-7B Llama Team (2024) and Qwen2.5-7B Qwen Team (2024). These models are evaluated in zero-shot and few-shot settings using a template adapted from Brinkmann et al. (2024), which incorporates the category, attribute, product profile, and detailed value normalization guidelines. We further fine-tune the LLMs using LoRA Hu et al. (2022) to improve performance.
RAG. We implement RAG baselines based on both Qwen2.5 and Llama3.1, using BGE embeddings Xiao et al. (2023) to retrieve the top-5 most similar product items from the training set. The LLM prompt is augmented with the profiles and structured attribute-value pair outputs of these retrieved examples. For all generation baselines, we apply greedy decoding to ensure reproducibility, and model outputs are formatted in JSON.
@@ 5 Results
5.1 Main Results
Table 4 presents the performance comparison between our retrieval-based method TACLR and classification- and generation-based baselines. On Xianyu-PAVI, TACLR achieves the highest F1 score of 86.2%, surpassing fine-tuned Llama3.1, which obtains 84.7%. Notably, TACLR excels in recall, achieving 87.1% compared to Llama3.1’s 82.7%. On WDC-PAVE, TACLR achieves the highest F1 Excl. score of 80.3%, which excludes measurement attributes requiring unit normalization reasoning. These results highlight TACLR’s effectiveness and robustness in addressing general PAVI across diverse datasets.
The classification-based baseline, BERT-CLS, shows the weakest performance on both datasets. It achieves an F1 score of 50.5% on Xianyu-PAVI, but its performance drops drastically to 20.5% on WDC-PAVE, underscoring the limitations of classification approaches in generalization. One contributing factor is the extreme label sparsity in this formulation: there are over 6.3M category-attribute-value labels, but the training set contains fewer than 3.6M instances, making it difficult for the model to learn a reliable classification head.
Among generation-based methods, performance improves steadily from zero-shot to few-shot, RAG, and fine-tuning on Xianyu-PAVI. Llama3.1 progresses from 35.7% (zero-shot) to 38.6% (few-shot), 47.6% (RAG), and 84.7% (fine-tuned), while Qwen2.5 improves from 48.4% to 51.4%, 63.2%, and 81.7%, respectively. A similar trend holds on WDC-PAVE, though fine-tuning is less effective due to limited supervision. In this setting, RAG outperforms few-shot prompting for both Qwen2.5 (74.2% vs. 72.2%) and Llama3.1 (77.2% vs. 75.0%), offering a viable alternative when labeled data is limited. While TACLR substantially outperforms RAG on Xianyu-PAVI (86.2% vs. 47.6%/63.2%), its advantage is narrower on WDC-PAVE, where it matches or slightly exceeds RAG in non-measurement attributes (F1 Excl.: 80.3% vs. 80.1%/78.3%). This likely stems from the limited size of WDC-PAVE, which constrains the effectiveness of supervised learning approaches, including BERT-CLS, fine-tuned LLMs, and TACLR.
Table 5: Inference efficiency comparison on Xianyu-PAVI (Throughput in samples/second).
@@ 5.1 Main Results
Table 6: F1 scores on product-attribute pairs with normalized vs. unnormalized and implicit values.
Method Normalized Unnorm. & Implicit
Llama3.1 83.2 79.4
Qwen2.5 82.6 78.6
TACLR 87.9 82.9
5.2 Inference Efficiency
Table 5 presents a comparison of inference efficiency across different PAVI paradigms under identical conditions, using an unoptimized PyTorch implementation on a single NVIDIA V100 GPU.
TACLR achieves a strong balance between model capacity and efficiency, processing each sample in 12.7 ms and achieving a throughput of 630 samples per second. In contrast, generation-based methods using Llama3.1 and Qwen2.5 exhibit substantially higher inference times (over 100 ms per sample) and lower throughput (below 100 samples per second), primarily due to the overhead of autoregressive decoding and reliance on large-scale models. The BERT-CLS classification baseline offers the highest raw efficiency (8.6 ms, 930 samples per second), but its inability to handle out-of-distribution values and limited generalization capacity reduce its practical applicability.
In summary, among these methods, TACLR provides the best trade-off between identification capability and inference efficiency, making it well-suited for scalable industrial deployment.
5.3 Analysis
@@
This section analyzes our method on the Xianyu-PAVI dataset, selected for its large scale and diverse attribute taxonomy. In contrast, the WDC-PAVE dataset is substantially smaller in both size and coverage, making it insufficient for robust evaluation of scalability and generalization.
@@
(c) Comparison of data domains.
@@
Robustness Evaluation on Diverse Values. To better assess model robustness under various conditions, we partition the test set into three subsets based on the nature of the ground truth: (1) normalized values, appear verbatim in the product profile; (2) unnormalized or implicit values, which either appear in a lexically varied form, or must be inferred; and (3) null values, where the attribute value is marked as unavailable. In our Xianyu-PAVI dataset, normalized values account for approximately 15.2%, unnormalized or implicit values for 27.2%, and null values for 57.6%.
Table 6 reports the F1 scores on the normalized and unnormalized/implicit subsets. TACLR achieves the highest performance on both subsets, with an F1 score of 87.9% on normalized values and 82.9% on unnormalized or implicit values. Compared to the fine-tuned LLM baselines, TACLR demonstrates stronger ability to recognize standardized values and better robustness to lexical variation and implicit inference.
@@
Impact of Taxonomy-Aware Negative Sampling. Figure 4 compares the proposed taxonomy-aware negative sampling (§3.3) with in-batch negative sampling across varying sample sizes. As the number of sampled values increases, the F1 score improves consistently, aligning with findings from Chen et al. (2020). Using in-batch sampling as the baseline, the model achieves an F1 score of 53.3% with a sample size of 128. In contrast, taxonomy-aware sampling yields substantial improvements, boosting the F1 score from 84.0% to 86.2% as the sample size grows from 16 to 128.
These results demonstrate that taxonomy-aware sampling provides more effective supervision, encouraging the model to distinguish between fine-grained, semantically similar values within the same category and attribute. In contrast, in-batch sampling often yields random negatives that are less relevant and frequently trivial, limiting the efficacy of contrastive learning for PAVI. For example, in Figure 1, negatives such as iPhone 12 Pro or iPhone 13 Pro Max, which belong to the model attribute under the phone category, provide more challenging and informative supervision than random values such as L for T-shirt size.
Comparison of Dynamic and Static Thresholds. Figure 5(a) compares our dynamic thresholding approach (§3.4) with static thresholds of 0.6, 0.65, and 0.7, selected via validation. The dynamic threshold achieves the highest F1 (86.2%), outperforming the static baselines (75.5%, 80.2%, and 78.2%). Static thresholds exhibit the typical precision–recall trade-off: higher thresholds increase precision (65.1%→84.5%) but reduce recall (90.0%→72.8%). In contrast, the dynamic threshold balances precision (85.4%) and recall (87.1%) without manual tuning. This gain stems from a scalable design: instead of relying on fixed, hand-tuned cutoffs per category-attribute pair, our method learns null value embeddings whose similarity scores act as adaptive thresholds.
@@
Performance Gains from Context-Rich Prompts. The influence of varying value prompt templates on the PAVI task is shown in Figure 5(b). Using only the value as a prompt achieves an F1 score of 83.2%. Adding category information raises the F1 score to 83.9%, while incorporating attribute information further improves it to 85.4%. The most comprehensive template, combining category, attribute, and value information (i.e., A {category} with {attribute} being {value}), achieves the highest F1 score of 86.2%. These results are consistent with prior work Radford et al. (2021), highlighting that context-rich prompts enhance the model’s discriminative performance.
Zero-Shot Generalization Across Data Domains. Figure 5(c) presents zero-shot transfer results on unseen categories and values. The in-domain split achieves an F1 score of 88.7%, while cross-category and cross-value splits decline to 80.2% and 78.2%, respectively, reflecting the challenges of adapting to evolving attribute taxonomies in out-of-distribution domains. Despite this, TACLR maintains a strong overall F1 of 86.2%, demonstrating robust generalization. The model’s generalization relies on the shared textual encoder’s semantic understanding and the retrieval-based approach’s capacity to leverage these embeddings for zero-shot matching. Such adaptability is critical for dynamic e-commerce platforms, where new products and attribute-value pairs continuously emerge, reducing the need for frequent retraining and lowering maintenance costs.
@@ 6 Conclusion
In this work, we present TACLR, a novel retrieval-based approach for PAVI. By formulating PAVI as an information retrieval problem, TACLR enables the inference of implicit values, generalization to OOD values, and the production of normalized outputs. Building on this framework, TACLR employs contrastive training with taxonomy-aware sampling and adaptive inference with dynamic thresholds to enhance retrieval performance and scalability.
Comprehensive experiments on proprietary and public datasets demonstrated TACLR’s superiority over classification- and generation-based baselines. Notably, TACLR achieved an F1 score of 86.2% on the large-scale Xianyu-PAVI dataset. Our efficiency analysis further highlighted its advantage, achieving significantly faster inference speeds than generation-based methods. Beyond these experimental results, TACLR has been successfully deployed on the real-world e-commerce platform Xianyu, processing millions of product listings daily and seamlessly adapting to dynamic attribute taxonomies, making it a practical solution for large-scale industrial applications.
@@ 7 Limitations
TACLR assumes access to a predefined attribute taxonomy, which serves as a foundation for accurate value identification. While this setup is realistic in many e-commerce platforms, where dedicated systems and teams maintain evolving taxonomies, it does require ongoing manual updates to incorporate new categories, attributes, and values. Automating product attribute mining remains an open area for future research Ghani et al. (2006); Zhang et al. (2022); Xu et al. (2023).
Another limitation is that TACLR currently operates on textual product profiles and does not incorporate multimodal information, such as images or videos. Multimodal inputs could provide complementary signals for attributes that are difficult to infer from text alone (e.g., color, material, or shape). Extending the method to support multimodal input could further improve its coverage and accuracy in practical applications.
A further limitation of TACLR, as a retrieval-based approach, is its difficulty in handling measurement attributes that require unit conversion or numerical reasoning. To address this, future work could explore hybrid methods that combine TACLR’s retrieval strengths with the reasoning ability of LLMs to improve performance in such cases.
@@ Appendix A Implementation Details
For both the BERT-CLS baseline and our TACLR method, we utilize pre-trained RoBERTa-base Liu et al. (2019); Cui et al. (2020) as the backbone. For TACLR, we augment the model with a linear projection head to map the embedding dimension to 256. For each product-attribute pair, we sample up to 128 values for the contrastive learning setup, including a null value, an optional positive value (if present for the attribute), and negative values sampled from the value set with the same category and attribute. In the case of the null value, we replace the value slot in the prompt template with null (e.g., A phone with capacity being null.).
The temperature parameter is fixed at 0.05. This choice is motivated by both prior research (e.g., SimCLR Chen et al. (2020) and MoCo He et al. (2020)) and empirical tuning on our validation set. The encoder is fine-tuned using the AdamW optimizer, with a batch size of 32, a learning rate of $2\times 10^{-5}$, and a maximum of 5 epochs.
For LLM fine-tuning, we employ LoRA Hu et al. (2022) for efficient adaptation. The core hyperparameters are as follows: 3 training epochs, batch size of 128, AdamW optimizer, maximum learning rate of $5\times 10^{-5}$, 1% warmup steps, cosine learning rate scheduler, LoRA rank of 8, LoRA alpha of 16, and LoRA dropout rate of 0.1.
All hyperparameters and model checkpoints are selected to maximize the F1 score on the validation set. For further details, please refer to our codebase: https://github.com/SuYindu/TACLR.
@@ Appendix B Deployment
The proposed TACLR has been successfully integrated into key functionalities of the e-commerce platform Xianyu, including product listing, search, recommendation, and price estimation. The system is designed to be highly scalable, efficiently processing millions of products daily.
During the product listing process, TACLR automatically identifies attribute–value pairs from user-provided titles and descriptions. This automation significantly reduces manual effort and errors, while enhancing the quality of structured product information.
For product search, the improved structured information directly supports more effective lexical retrieval and enables structured search, where items are filtered based on attribute values. This leads to more accurate matching with user queries. In addition, the enriched product features enhance the quality of personalized recommendations.
In the context of price estimation, TACLR identifies key attribute values that influence pricing, enabling more accurate price predictions. This functionality provides sellers and buyers with reliable, market-aligned information in the context of second-hand transactions.